OpenStudy (anonymous):

help me factor out this polynomial 1/4x^2+x+1 please.

OpenStudy (ikram002p):

multibly by 4 x^2+4x+4 =(x+2)^2 = (x+2) (x+2) :3

OpenStudy (anonymous):

thank you so much !!!!!!

OpenStudy (ikram002p):

ur very wlc